OpenStudy (anonymous):

If line segment KN is a median of ΔKLM, find KL. C.9 or D.11 ?

OpenStudy (anonymous):



OpenStudy (e.mccormick):

Do you know the definition of the median?

OpenStudy (anonymous):

the line in the middle right? ..i just learned these terms yesterday. :/ so im not sure.

OpenStudy (e.mccormick):

Middle is close. It connects a vertex to the opposite midpoint. The midpoint means it is half way between. So, if N is half way between M and L, do you see any mathematical relationships you can use?

OpenStudy (anonymous):

2x+3+3x-7 ??

OpenStudy (e.mccormick):

Forget about \(\overline{KL}\) for now. Start with just \(\overline{ML}\). Because N is half way between M and L, what does that mean about the mause of \(\overline{MN}\) and \(\overline{NL}\)?

OpenStudy (e.mccormick):

measure.

OpenStudy (anonymous):

ok, it means its halfway across the base. what do you mean by measure??

OpenStudy (anonymous):

Do you mean 3(5)-7=8 Like that?

OpenStudy (e.mccormick):

It rearranged my post order. That was odd. Measure, to find the lenght. Lets get away from even that for a moment. If I take some line segment, any line segment, and I cut it in half, what one fact would I know about how long both halfs were?

OpenStudy (anonymous):

they would be equal halfs.

OpenStudy (e.mccormick):

The big word there being equal. OK. So, you have been told you have a median. By definition, that median has a point that cuts the side in half. That means the segements it is cut into are equal. If they are equal, what equation can you make from the drawing?

OpenStudy (anonymous):

3x-7

OpenStudy (e.mccormick):

\(3x-7=?\;\leftarrow\) an equation has an equals sign. So what is the other side of the equation?

OpenStudy (anonymous):

3x-7=5 or 3x-7=3x-7

OpenStudy (e.mccormick):

Yes, 3x-7=3x-7, but the one I was getting at was: \(\overline{NL}=\overline{MN}\implies 3x-7=5\) See, knowing the terms lets you set up that equation. That equation lets you solve for x. Once you know x you can solve any other part of that triangle with an x in it. =)

OpenStudy (anonymous):

ok how do i get x?

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

x=4

OpenStudy (e.mccormick):

Yes. That is it at that point.

OpenStudy (anonymous):

my answer 11. thanks!

OpenStudy (e.mccormick):

Now that you have x, you can use it to solve.... hehe. I think you got it!